A spring has an unstretched length of 15 cm . When an 85 g ball is hung from it, the length increases by 3.0 cm . Then the ball is pulled down another 3.0 cm and released.

Part A) What is the spring constant of the spring? Express your answer in newtons per meter.

Part B )What is the period of the oscillation? Express your answer in seconds.
Answer #1

A)

Force is given by,

F = mg = kx

Spring constant, k = mg/x = (0.085 x 9.8)/0.03 = 27.77 N/m

B)

Time, T = 2 pi sqrt(m/k) = 2 pi sqrt(0.085/27.77) = 0.35 sec

Comment in case any doubt please rate my answer ...